Ticket Number: HEA-NOC/20031124-7 Ticket Status: OPEN
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20031124 18:14 UTC Problem Start: 20031127 10:00 UTC
Ticket Closed: Problem End: 20031127 12:00 UTC
On Thursday, 27/11/03, there will be a short outage on ftp.heanet.ie to
facilitate a physical move to a new cabinet, a kernel upgrade, and the
migration of the storage array to an NFS backend server. This outage is
expected to last in the region of 30 minutes within the maintainence
window of 2 hours.
All access to ftp.heanet.ie will be limited during the outage.
20031124 18:06 CMC Opening ticket.
This ticket can be monitored at http://www.hea.net/tickets/20031124-7